[figure: A; B; B; B; B; B; B; B; B; B; B; B; B; B; B; B; B; B; Point 1; Point 2; Point 3]

Figure 1. My first picture

This document will show some of the ways “pictures” can be included in a L^{A}TEX2_{ϵ} document.

Comments

• The length \unitlength should be relatively small since all coordinates (including the picture size) must be integers.

Top/Right (no frame)

Top/Left Bottom/Right

Bottom/Left

[figure] [figure] [figure]

[figure]

Figure 2. And another